# Firstname1 Lastname1    Firstname2 Lastname2    Firstname3 Lastname3    Firstname4 Lastname4    Firstname5 Lastname5    Firstname6 Lastname6    Firstname7 Lastname7    Firstname8 Lastname8    Firstname8 Lastname8

### v1

@@ 5 Scaling Up the Modulus
Figure 8: Visualization of byte tokenization and abacus embedding. Abacus embedding 1 is shared by all the bytes within the integer, while Abacus embedding 2 varies within the byte but is shared by all integers.
@@
In this section, we investigate Transformer training upon scaling up the modulus of LCG. For this purpose, we make the following modifications to our model:
@@
Byte Representation Inspired by the emergent radix representations observed in the previous section and seeking to avoid massive dictionary sizes for large $m$, we implement a byte-level tokenization scheme. Each integer is decomposed into a sequence of bytes, beginning with the least significant byte. The vocabulary has a size of $256$ for any $m$.
@@
Abacus Embeddings We encode positional information using a variant of the Abacus embedding (mcleish2024abacus), as a sum of two learnable vectors. One vector encodes the position of the integer within the sequence, while the other encodes the position of each byte within the integer. In Figure 8, we show a visual representation.
@@
Fixed Modulus For FM training, we train a 2-layer model with $d_{\mathrm{model}}=1024$ and a vocabulary size of $256$. We select training sequences via the Hull-Dobell theorem, setting $n_{a}=n_{c}=1024$ (See Section F.1 for training details). For the test dataset, we choose $a$ and $c$ that differ from those in the training set. The quality of an LCG largely depends on its multiplier, traditionally evaluated via the spectral test. In Figure 9(a), we test our model on both spectrally optimal Steele multipliers (steele2021) and arbitrary multipliers for $m=2^{32}$. While achieving $100\%$ test accuracy with equal in-context sequence lengths, the model performs consistently worse on Steele-generated sequences compared to those from arbitrary multipliers.
@@
In Figure 9(b), a log-log plot reveals that the number of in-context examples needed for $100\%$ test accuracy scales sublinearly with modulus $m$ as $m^{\alpha}$, where $\alpha\approx 1/4$.
@@
Figure 9: (a) Average test accuracy ($m=2^{32}$) vs number of in-context examples. (b) The number of in-context examples required to achieve 100% test accuracy (minimum of 5 runs). (See Appendix E for details)
@@
Generalization to unseen modulus For UM case, we train a 6-layer Transformer on a dataset with $n_{m}=32,768,n_{a}=128,n_{c}=1$ ($655361024<m_{\mathrm{train}}<$ 65536), with 512 numbers in each sequence. As before, we use unseen values of $m_{test}$ and $a,c$ for the test dataset – with a focus on $m_{test}$ of the form $2^{k}$ and $3^{k}$. In Figure 10, we show how the number of in-context sequence elements is required to reach $60\%$ test accuracy scales with $m$. The averaged test accuracy of each number in the sequence is shown in Section F.2. Test performance is influenced by the tokenization base, since the tokenization base highlights the periodic structure of LCGs making it more apparent and easier for the model to leverage during training and prediction. In Figure 10 we observe that $m=2^{k}$ ($m=3^{k}$) sequences offer favourable scaling when the tokenization base is $256=2^{8}$ ($243=3^{5}$).
@@
Figure 10: In-context eample scaling for UM: The number of in-context sequence elements required to achieve 60% test accuracy shows a sublinear dependence on the modulus $m$. The scaling depends on the compatibility between $m$ and tokenization. (a) base-$2^{8}$ tokenization, (b) base-$3^{5}$ tokenization.
6 Discussion
@@
We have investigated Transformer training on LCG sequences, focusing on fixed modulus training as well as generalization to unseen moduli. In both cases, we have uncovered the algorithm used by the model to solve these tasks, and highlighted the model components that implement the steps of the algorithm. We have found that the model finds and utilizes prime factorizations of $m$ and mixed-radix representations of numbers to simplify the sequences and make predictions. We have provided the modified training recipe for scaling up the modulus in both FM and UM settings, and shown their scaling behaviors.
Limitations and future work: While we scale the FM setting to $m=2^{32}$ the We only scale UM setting up to $m=2^{16}$ in this work. We also leave the exploration of PRNGs that are built upon LCG, such as PCG and TCG for future works.
@@
Appendix A Fixed Modules Training Results
@@
For the FM case, we find that a single attention head in one layer is sufficient to solve the task. In Figure 11, we present the model’s training loss and performance across training steps for $m=2048$. Notably, panels (c, f) reveal a significant disparity between training and test loss, indicating a grokking transition during the training process. Similarly, we plot in Figure 12 for similar curves for $m=776$, where all curves are qualitatively the same.
@@
Figure 11: Test accuracy and train/test loss for $m=2048=2^{11}$, depth=1. (a,b,c) $n_{heads}=1$ (d,e,f) $n_{heads}=4$.
@@
Figure 12: Test accuracy and train/test loss for $m=7776=2^{5}\cdot 3^{5}$, depth=1. (a,b,c) $n_{heads}=1$ (d,e,f) $n_{heads}=4$.
Appendix B Critical Depth for the Unseen Modulus Task
@@
Figure 13: Test accuracy heatmaps of
@@
This section analyzes the depth and embedding dimension requirements for successfully training a Transformer on the unseen modulus task. We generate $N=4\times 10^{5}$ training examples covering at least $m_{\text{eval}}/4$ moduli in the training dataset.
@@
We varied depths $\in\{2,3,4,6,8\}$ and embedding dimensions in $d_{\text{model}}\in\{512,768,1024,1280\}$, with the head dimension fixed to $d_{\text{head}}=128$. For each depth and width, we scanned learning rates $\eta\in\{3\times 10^{-5},1\times 10^{-4},3\times 10^{-4},1\times 10^{-3}\}$ and weight decay strengths $\lambda\in\{0.01,0.1,1.0,3.0\}$ to identify the optimal hyperparameters. We report that the optimal learning rate and weight decay strength heavily vary with depth and embedding dimension and training dataset. For $m_{\text{eval}}=\{1024,4096\}$, the models were trained for $T=100,000$, while for $m_{\text{eval}}=16,384$, the models required a longer training for $T=200,000$ steps. In both cases, we used a batch size of $256$ and a linear learning rate warmup over $4,096$ steps using AdamW with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.99$.
@@
Appendix C Training Time Interpratibility
@@
Appendix D More Interpretability
D.1 Fixed Modulus
In this subsection, we show extra results on the model’s behavior…
Figure 15: Test accuracy for $m=2048=2^{11}$, depth=1, $n_{heads}=1,d_{model}=768$. (a) Test accuracy averaged over $a,c$ and initial seeds. (b) Multiplication of per-digit test accuracies – matches exactly with the average test accuracy. (c) Per-digit accuracy in binary representation. (d) Per-digit accuracy of an ablated model: For each query, mask out all attention weights except for the top 2, which are $2^{k}$ and $2^{k-1}$ distance backwards in the context
Figure 16: Test accuracy for $m=2048=2^{11}$, depth=1, $n_{heads}=4,d_{model}=768$. (a) Test accuracy averaged over $a,c$ and initial seeds. (b) Multiplication of per-digit test accuracies – matches exactly with the average test accuracy. (c) Per-digit accuracy in mixed radix representation.

### v2

@@ 5 Scaling Up the Modulus
In this section, we investigate training upon scaling up the LCG modulus, with the following modifications:
@@
Base-b tokenization: To avoid massive dictionary sizes for large $m$, we implement a base-$b$ tokenization scheme. Each integer is decomposed into a sequence of base-$b$ digits, beginning with the least significant digit; resulting in a vocabulary size of $b$ for any $m$ (for details see Section I.1). Based on the discussion in Section 4.1, it is beneficial to choose $b$ such that $\mathrm{gcd}(b,m)=b$.
@@
Abacus Embeddings: We encode positional information using a variant of the Abacus embedding (McLeish et al., 2024), as a sum of two learnable vectors. One vector encodes the position of the integer within the sequence, while the other encodes the position of each digit within the integer. (For details, see Section I.2).
Figure 11: Visualization of base-$b$ tokenization and abacus embeddings. Abacus embedding 1 is shared by all the digits within the integer, while Abacus embedding 2 varies within the digit but is shared by all integers.
5.1 Fixed Modulus
@@
For each modulus $m=2^{k}$, where k is an integer in the range $16\leq k\leq 32$, we train a 2-layer model with $d_{\mathrm{model}}=1024$ and a vocabulary size of $256$. We select training sequences via the Hull-Dobell theorem, setting $n_{a}=n_{c}=1024$ (See Section I.3 for training details). For the test dataset, we choose 512 values of $a$ and 64 values of $c$ that differ from those in the training set.
The quality of an LCG largely depends on its multiplier, traditionally evaluated via the spectral test (Knuth, 1997). In Figure 12(a), we test our model on both spectrally optimal Steele multipliers (Steele & Vigna, 2021) and arbitrary multipliers for $m=2^{32}$. While achieving $100\%$ test accuracy with equal in-context sequence lengths, the model performs consistently worse on Steele-generated sequences compared to those from arbitrary multipliers.
@@
In Figure 12(b), a log-log plot reveals that the number of in-context sequence elements needed for $100\%$ test accuracy scales sublinearly with modulus $m$ as $m^{\gamma}$, where $\gamma\approx 1/4$.
@@
Figure 12: FM: (a) Average test accuracy ($m=2^{32}$) vs number of in-context sequence elements. (b) The number of in-context sequence elements required to achieve 100% test accuracy (minimum of 5 runs). (See Section I.3 for details)
5.2 Unseen modulus
@@
For the UM case, we train a 6-layer Transformer on a dataset with $n_{m}=32{,}768$, $n_{a}=128$, $n_{c}=1$, with $5361024<m_{\mathrm{train}}<65{,}$ 536. Sequences are length 512, each integer tokenized as two bytes (context length 1023). As before, test data uses unseen $m_{\mathrm{test}}$ and $(a,c)$, focusing on $m_{\mathrm{test}}=2^{k},,3^{k}$. Figure 13 shows that the number of in-context sequence elements needed to reach $60\%$ test accuracy scales as $m_{\mathrm{test}}^{\gamma}$ ($0.24\leq\gamma\leq 0.33$). The averaged test accuracy of each number in the sequence is shown in Section I.4. Test performance is influenced by the tokenization base, since the tokenization base highlights the periodic structure of LCGs making it more apparent and easier for the model to leverage during training and prediction. To confirm this, we train a model with tokenization base 243. In Figure 13, $m_{\mathrm{test}}=2^{k}$ ($3^{k}$) sequences scale better when the tokenization base is $256=2^{8}$ ($243=3^{5}$).
@@
Figure 13: UM: The number of in-context sequence elements needed for $60\%$ test accuracy grows sublinearly with modulus $m$, depending on compatibility between $m$ and tokenization. (a) Base-$2^{8}$ tokenization; (b) base-$3^{5}$ tokenization.
6 Conclusion
@@
We have investigated Transformer training on LCG sequences, focusing on fixed modulus training as well as generalization to unseen moduli. In both cases, we have uncovered the algorithm used by the model to solve these tasks and highlighted the model components that implement the steps of the algorithm. We have found that the model finds and utilizes prime factorizations of $m$ and RNS representations of numbers to simplify the sequences and make predictions. We have provided the modified training recipe for scaling up the modulus in both FM and UM settings, and shown their scaling behaviors.
Limitations and future work: The results of this paper were limited to scales $m\leq 2^{32}$. It would be interesting to test our results on much larger moduli as well. We leave the exploration of PRNGs that are built upon LCGs, such as PCGs and truncated LCGs for future works. It would also be interesting to make the training even more unbiased, by training on general classes of arithmetic sequences.
Acknowledgements
M.B. thanks Carl Miller for discussions on PRNGs. This work is supported by NSF DMR-2345644 (D.S.K., T.T., and M.B.), and by an NSF CAREER award DMR-2045181 (T.H. and D.D.). The authors acknowledge the University of Maryland supercomputing resources (http://hpcc.umd.edu) made available for conducting the research reported in this paper.
@@ Acknowledgements
Appendix A Experimental Details
This section provides further details about model architecture, dataset construction, and optimization.
A.1 Dataset Construction
Fixed Modulus (FM): Given a modulus $m$, we apply the Hull-Dobell Theorem to determine the possible values of ($a,c$) that maximize the period. We then randomly select $64$ values of $a$ and $c$ to generate the test dataset. To generate the training dataset, we exlude these test choices of ($a,c$) and uniformly sample $N=100,000$ LCG sequences of length $L$ (context length) with $n_{a}$ values of multipliers and $n_{c}$ values of increments. For each set of parameters ($a,c$), we sample an LCG sequence with a randomly selected initial seed $x_{0}$. Note that the training dataset includes sequences with varying periods, while the test data only contains sequences that maximize the period.
Figure 14: The impact of training dataset parameters ($n_{m},n_{a},n_{c}$) on unseen modulus task performance.
Generalization to Unseen Modulus (UM): For the test dataset, we first select a set of test moduli $M_{\text{test}}=\{m_{\text{test}}\}$ that would be reserved exclusively for evaluation. For each test modulus $m_{\text{test}}\in M_{\text{test}}$, we apply the Hull-Dobell Theorem to determine the values of ($a,c$) that maximize the period. We then randomly select $64$ values of $a$ and $c$ to generate the test dataset. These $64$ ($a,c$) values are not considered while generating the training dataset.
For the training dataset generation, we sample the $n_{m}$ modulus values from the range $[L,\lfloor 1.2\max(M_{\text{test}})\rfloor]$ while excluding all the values in $M_{\text{test}}$. For each modulus value $m$, we uniformly select $n_{a}$ multipliers ($0<a<m$) and $n_{c}$ increments ($0\leq c<m$), excluding the ones reserved for testing. For each parameter ($a,c,m$), we generate a sequence of length $L$ using a randomly selected initial seed $x_{0}$. This results in a total of $N=n_{m}\times n_{a}\times n_{c}$ training sequences. We report that $N=400,000$ served sufficient from the modulus values considered in this work.
Next, we examine the effect of training dataset composition ($n_{m},n_{a}$, $n_{c}$) on the performance. Figure 14 shows the test accuracy primarily depends on $n_{p}$ and marginally on $n_{a}$ and $n_{c}$. Furthermore, we found that $n_{m}\gtrsim m_{\text{test}}/4$ yields good generalization performance (result not shown here). Based on this relationship and our target total number of training examples $N$, we sample $n_{a}=n_{c}=\sqrt{\frac{N}{m_{\text{test}}/4}}$ values of multipliers and increments. Unless explicitly specified, we use these parameter settings as the default configuration for all experiments.
A.2 Model Architecture Details
We consider GPT-style Transformers (Radford et al., 2019) with learnable positional embeddings and weight tying (Press & Wolf, 2017). The model architecture is characterized by the number of blocks (depth), embedding dimension ($d_{\text{model}}$), and number of attention heads ($n_{\text{heads}}$). For most experiments, we use GELU activations, except in Section 4, where we use ReLU activations for better interpretability.
A.3 Training Details
Figure 15: Heatmap of test accuracy of a $6$ layer Transformer with learning rate and weight decay as the axes. As the modulus is increased from $1024$ to $16,384$ the range of hyperparameters resulting in reasonable accuracy becomes narrow.
We train the models with Cross-entropy loss using AdamW optimizer (Loshchilov & Hutter, 2019) with momentum hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.99$. We implement a linear learning rate warmup over the first $2048$ steps with an initial learning rate of zero and the target learning rate $\eta$. By default, all experiments employ a batch size of $256$. Weight decay is only applied to non-bias parameters.
In the unseen modulus case, we observed that both the optimal target learning rate $\eta$ and weight decay strength $\lambda$ are highly sensitive to minute changes in training dataset properties (modulus $m_{\text{test}}$, number of LCG parameters $n_{m}$, $n_{a}$, $n_{c}$ and total number of examples $N$) and architectural changes (depth and embedding dimension). To determine the optimal hyperpamraters, we scan the learning rates $\eta\in\{3e\text{-}05,1e\text{-}04,3e\text{-}04,1e\text{-}03\}$ and weight decay strengths $\lambda\in\{0.01,0.1,1.0,3.0\}$.
A.4 Training Cost
In Figure 12, the training of the $m=2^{32}$ model was conducted using four NVIDIA A100 GPUs, requiring a total of 21.82 hours. The $m=2^{16}$ model completed training in 4.83 hours under the same hardware setup. Despite having the same model size, the increased context length in the $m=2^{32}$ model led to a significantly higher computational cost.
In Figure 13, both models were trained on a single NVIDIA H100 GPU for 22 hours.
A.5 Hyperparameter Space Shrinking with increasing modulus
We also report a surprising phenomenon in the unseen modulus case, which we refer to. as the ‘hyperparameter space shrinking.’ As we increase the test modulus $m_{\text{test}}$ while keeping the model architecture fixed, we observe that the optimal learning rate ($\eta$) and weight decay strength ($\lambda$) shift significantly. Moreover, the range of these hyperparameter values that yield reasonable performance becomes increasingly narrow. Figure 15 shows this result for a $6$ layer Transformer.
Appendix B Fixed Modules Training Results
@@
For the FM case, we find that a single attention head in one layer is sufficient to solve the task when the modulus is not a prime number. In Figure 16, we present the model’s training loss and performance across training steps for $m=2048$. Notably, panels (c, f) reveal a significant disparity between training and test loss, indicating a grokking transition during the training process. Similarly, we plot in Figure 17 for similar curves for $m=7776$, where all curves are qualitatively the same.
@@
Figure 16: Test accuracy and train/test loss for $m=2048=2^{11}$, depth=1. (a,b,c) $n_{\mathrm{heads}}=1$ (d,e,f) $n_{\mathrm{heads}}=4$.
@@
Figure 17: Test accuracy and train/test loss for $m=7776=2^{5}\cdot 3^{5}$, depth=1. (a,b,c) $n_{\mathrm{heads}}=1$ (d,e,f) $n_{\mathrm{heads}}=4$.
Appendix C Prime Moduli
In the FM setting, when $m$ is a prime number, the task becomes much harder. Since there are no digit-wise periodic patterns, the model cannot perform the algorithm described in Section 4.2. In Figure 18, we trained two identical models to learn $m=2039$ and $m=2048$, and we observe that the task with $m=2039$ cannot be learned within the same number of training steps. Note that to rule out potential constraints from model capability, we used depth 2 models (as oppposed to depth 1 in the main text) in both cases.
(a) Test Accuracy
(b) Training Accuracy
(c) Training Loss
Figure 18: FM: Comparison between $m=2039$ (prime) and $m=2048$ (power-of-two), with both models trained for 50,000 steps. Each model has depth 2 and $d_{\text{model}}=1024$. The test set consists of sequences with the same periods, while the training set includes arbitrary multipliers not present in the test set.
In the UM setting, the model exhibits similar test performance on sequences with prime $m$ and sequences with $m$ as a power of two, as shown in Figure 19. We hypothesize that training on a diverse set of moduli helps the model rely less on the digit patterns and instead focus on more generalizable structure.
Figure 19: UM: Test accuracy comparison between $m=2039$ (prime), $m=1801$ (prime) and $m=2048$ (power-of-two). The model was trained for 100,000 steps on a dataset consisting of 262,144 sequences with 512 distinct training moduli not present in the test set.
Appendix D Critical Depth for the Unseen Modulus Task
@@
Figure 20: Test accuracy heatmaps of with depth and embedding dimensions as the two axes.
@@
This section analyzes the depth and embedding dimension requirements for successfully training a Transformer on the unseen modulus task. The experimental details are the same as described in Appendix A.
@@
We varied depths $\in\{2,3,4,6,8\}$ and embedding dimensions in $d_{\text{model}}\in\{512,768,1024,1280\}$, with the head dimension fixed to $d_{\text{head}}=128$. For each depth and width, we scanned learning rates $\eta\in\{3e\text{-}05,1e\text{-}04,3e\text{-}04,1e\text{-}03\}$ and weight decay strengths $\lambda\in\{0.01,0.1,1.0,3.0\}$ to identify the optimal hyperparameters. We report that the optimal learning rate and weight decay strength heavily vary with depth, embedding dimension, and training dataset. For $m_{\text{eval}}=\{1024,4096\}$, the models were trained for $T=100,000$ steps, while for $m_{\text{eval}}=16,384$, the models required a longer training for $T=200,000$ steps.
@@
Appendix E Training Time Interpretability
@@
Appendix F LCG Properties
F.1 The period of the lower $k$-th bit when $m$ is power of 2
In this section, we show that for a sequence of period $\mathcal{T}_{m}=m=2^{K}$, the $k$-th lowest digit has a period of $2^{k}$ along the sequence. Consider an LCG sequence:
$x_{t+1}=(ax_{t}+c)\mod m,$ (6)
where $m$ is a power of 2, $m$ and $c$ are coprime and $a-1$ is divisible by 4. The period of sequence $x_{t}$ is m (Hull & Dobell, 1962). The lower $k$-th bits of $x_{t}$ is given by:
$b_{t,k}=\frac{z_{t,k}-z_{t,k-1}}{2^{k-1}},$ (7)
where $z_{t,k}=x_{t}\mod 2^{k}$. Therefore, if $z_{t,k}$ has a period of $2^{k}$, then the lower $k$ bits also have a period of $2^{k}$. Below, we show that $z_{t,k}$ has a period of $2^{k}$.
For an integer $M_{t}$, we can can re-write $x_{t}$ as:
$x_{t}$ $=z_{t,k}+M_{t}2^{k}.$ (8)
Next, we substitute Equation 6 into the definition of $z_{t+1,k}$:
$z_{t+1,k}$ $=x_{t+1}\mod 2^{k},$
$=[(ax_{t}+c)\mod m]\mod 2^{k}..$ (9)
As $m$ is divisble by $2^{k}$, this simplifies to:
$z_{t+1,k}$ $=(az_{t,k}+c)\mod 2^{k},$
$=(az_{t,k}+aM_{t}2^{k}+c)\mod 2^{k},$
$=(az_{t,k}+c)\mod 2^{k}.$ (10)
Therefore, $z_{t,k}$ follows its own LCG recurrence with the same $a$ and $c$ but with a reduced modulus $2^{k}$. Because $2^{k}$ and $c$ are coprime and $a-1$ is divisible by 4, the period of $z_{n}$ is $2^{k}$. Thus, the period of the lower $k$ bits is $2^{k}$.
Since $z_{n,k}$ has period $2^{k}$ and $z_{n,k-1}$ has period $2^{k-1}$, the period of $b_{n,k}$ is $2^{k}$.
F.2 Derivation of Equation 4
We now derive Equation 4 of the main text:
$x_{t}\;\mathrm{mod}\;p_{i}^{w_{q}}=\alpha_{i,0,t}\,p_{i}^{0}+\alpha_{i,1,t}\,p_{i}^{1}+\cdots+\alpha_{i,w_{i}-1,t}\,p_{i}^{w_{i}-1},$
where $\alpha_{i,w,t}\in\{0,1,\dots,p_{i}-1\}$ are base-$p_{i}$ digits. When the period of $x_{t}$ is m, each digit $\alpha_{j,w,t}$ has a period of $p_{i}^{w}$.
Consider an LCG sequence:
$x_{t+1}=(ax_{t}+c)\mod m,$ (11)
where $m$ has a prime factorization $m=p_{1}^{w_{1}}p_{2}^{w_{2}}\cdots p_{q}^{w_{q}}$, $m$ and $c$ are coprime, $a-1$ is divisible by all prime factors of $m$ and $a-1$ is divisible by 4 if $m$ is divisible by 4 (Hull & Dobell, 1962).
Consider the residulas $R_{i,t}=x_{t}\mod p^{w_{i}}_{i}$, where $i\in\{1\dots q\}$. We have:
$x_{t}=R_{i,t}+M_{i,t}p^{w_{i}}_{i},$ (12)
where $M_{i,t}$ is an integer.
Substituting Equations 11 and 12 into the definition of $R_{i,t+1}$:
$R_{i,t+1}$ $=x_{t+1}\mod p^{w_{i}}_{i},$
$=[(ax_{t}+c)\mod m]\mod p^{w_{i}}_{i},$
$=(ax_{t}+c)\mod p^{w_{i}}_{i},$
$=(aR_{i,t}+aM_{i,t}p^{w_{i}}_{i}+c)\mod p^{w_{i}}_{i},$
$=(aR_{i,t}+c)\mod p^{w_{i}}_{i}.$ (13)
Next we consider $z_{i,k,t}$, which is the lower $k$ base-$p_{i}$ digits of $R_{i,t}$:
$z_{i,k,t}=R_{i,t}\mod p_{i}^{k}.$ (14)
Similarly, the recurrence of $z_{i,k,t}$ can be simplied as:
$z_{i,k,t+1}$ $=R_{i,t+1}\mod p_{i}^{k},$
$=[(aR_{i,t}+c)\mod p_{i}^{w_{i}}]\mod p_{i}^{k},$
$=(aR_{i,t}+c)\mod p_{i}^{k},$
$=(az_{i,k,t}+c)\mod p_{i}^{k}.$ (15)
$z_{i,k,t}$ follows an LCG recurrence with the same $a$ and $c$ and a reduced modulus $p^{k}_{i}$. Because $p^{k}_{i}$ and $c$ are coprime and $a-1$ is divisible by $p_{i}$, the period of $z_{i,k,t}$ is $p^{k}_{i}$.
$\alpha_{i,k,t}$, which is the lower $k$-th base-$p_{i}$ digits of $y_{i,t}$, can be written as:
$\alpha_{i,k,t}=\frac{z_{i,k,t}-z_{i,k-1,t}}{p_{i}^{k-1}}.$ (16)
Since $z_{i,k,t}$ has period $p^{k}_{i}$, the period of $\alpha_{i,k,t}$ is $p^{k}_{i}$.
Appendix G Fixed Modulus Interpretability
In this subsection, we show additional results on the model’s behavior on the FM task.
Figure 22: Test accuracy vs token positions for $m=2048=2^{11}$, depth=1, various values of $a,c$. (a,b) $n_{\mathrm{heads}}=1$ (c,d) $n_{\mathrm{heads}}=4$. The accuracies for different $a$ and $c$ are exactly on top of each other.
Figure 23: Test accuracy for $m=2048=2^{11}$, depth=1, $n_{\mathrm{heads}}\in\{1,4\},d_{\mathrm{model}}=768$. (a,d) Test accuracy averaged over $a,c$ and initial seeds. (b,e) Multiplication of per-digit test accuracies – matches exactly with the average test accuracy. (c,f) Per-digit accuracy in binary representation.